6,x,y,16 solve for value of x
Answer by richwmiller(17219)

an = a + (n - 1)d
16 = 6 + (3)d
d=10/3
6,x,y,16
18/3,28/3,38/3,48/3
6,28/3,38/3,16
x=28/3